# Dynamical Similarity in Multisymplectic Field Theory

### v3

@@ V Field Theory Classification
VI Conclusions and Outlook
The use of multisymplectic geometry to provide a manifestly covariant description of classical field theory has recently assumed a more prominent role within the literature, due in part to the success of standard symplectic geometry, which continues to provide insightful connections between classical mechanics and differential geometry [92]. The finite dimensionality of the multisymplectic phase space is highly attractive, and constitutes a notable advantage over the canonical field-theoretic framework. The formalism of multicontact geometry, by contrast, is far less developed; action-dependent field theories provide a way to describe systems which exhibit non-conservative effects, and are thus encountered somewhat less frequently.
Following a brief review of the mathematical preliminaries required to treat classical field theories in a manifestly covariant manner, we introduced the notion of dynamical similarity, and demonstrated that the process by which a redundant ‘scaling’ degree of freedom may be eliminated has a well-defined generalisation to a field-theoretic context. In particular, we showed that a multisymplectic Lagrangian (or Hamiltonian) system admitting a dynamical similarity could be reduced to an action-dependent theory, defined on a lower-dimensional space. From a geometrical perspective, this symmetry reduction process had a simple yet elegant interpretation: the removal of a single degree of freedom $\rho$ leaves each of the velocities $\rho_{\mu}$ with no corresponding field variable; in the reduced theory, each of these unpaired velocities assumes the role of a component of the action density $s^{\mu}$, which, together, span a Reeb distribution, making the resulting theory multicontact.
@@ VI Conclusions and Outlook
Through use of a pair of simple examples, we found that, in general, a theory of $N$ non-interacting massive scalar fields, which possesses a scaling symmetry, is dynamically-equivalent to a configuration of (N1)(N-1) massless fields moving in a non-zero potential, set by the value of the original field mass, together with a completely independent frictional component, which we interpret as arising due to the empirical inaccessibility of the scaling variable. Furthermore, it was observed that any coupling between the scalar fields of the original multisymplectic theory destroys this separation into independent pieces, and introduces terms which mix the frictional degrees of freedom with those of the remaining scalar fields.
The results presented throughout constitute a foundational framework, within which a number of interesting areas of future work may be explored. For example, we have focused exclusively on regular Lagrangian systems; the regularity of our field theory guarantees that symmetry reduced space admits a well-defined multicontact structure. The vast majority of theories of physical interest possess gauge symmetries, which give rise to singular Lagrangian systems. In this case, we must develop a formalism adapted to the treatment of such theories, and, in particular, it will be of great interest to study the interaction between gauge fixing and contact reduction. At present, it remains unclear as to whether the process of pre-symplectic reduction, used to define the physical phase space, commutes with the elimination of a scaling degree of freedom [93]. Such a framework will also provide an arena within which to study General Relativity; it has been mentioned several times that, upon making a conformal decomposition, the Einstein-Hilbert Lagrangian possesses a dynamical similarity. Having developed a sound understanding of the interaction between gauge and scaling symmetries, it will be of great interest to examine the solution space of contact-reduced General Relativity, particularly at points where our conventional description becomes singular and thus ill-defined. It is known, for instance, that within the cosmological sector, both FLRW and certain classes of Bianchi models may be extended beyond the initial singularity [94, 95].
@@
In addition to the treatment of singular field theories, it will also be of interest to consider the behaviour of systems possessing fermionic degrees of freedom. Some progress has been made in this area; for example a De-Donder Weyl Hamiltonian description of a simple two-dimensional supersymmetric theory has been developed in [96]. Moreover, generalised Kähler geometries have been used to construct covariant Hamiltonian formulations of non-linear sigma models [97], and so an extension of our formalism to encompass anticommuting variables is highly desirable.
@@
Finally, we have mentioned several times throughout the advantages presented by the covariant De-Donder Weyl formulation of classical field theories. Whilst manifest covariance and finite-dimensional phase spaces are indeed attractive features of this approach, when contrasted with the canonical formalism, it cannot be overlooked that the latter provides a clear and rigorous path to quantisation, whereas the former does not. In light of this, of particular importance is the development of a standardised framework within which the restricted multimomentum bundle $J^{1}E^{*}$ may be quantised, in much the same way that geometric quantisation offers a scheme for performing such calculations on the tangent bundle.

### v4

@@ VII Field Theory Classification
VIII Singular Lagrangians and Further Applications
Before concluding, we make a number of remarks concerning singular field theories and further applicability of our formalism. Most field theories within contemporary theoretical physics possess gauge symmetries, and so are described by singular Lagrangians. The framework developed throughout is, at present, valid only for regular systems. Whilst the procedure for treating theories with gauge symmetries is, in principle, analogous to that presented here, there still remain several technical details that are unclear. For instance, given a singular multisymplectic Lagrangian $L$, one has two conceivable ways to proceed: direct contact reduction of $L$ to obtain a degenerate Herglotz Lagrangian, followed by multiphase space restriction via the geometrical algorithm [26, 25]. Alternatively, the constraint algorithm could be applied at the multisymplectic level, finding the maximal subset of $J^{1}E$ upon which solutions of the equations of motion exist, before seeking a scaling symmetry within the constrained theory. More work is required to ascertain the commutative relationship between these two approaches. Additionally, the geometrical interpretation of the reduced space, discussed at the end of section (VI), must be suitably modified, when we are confined to work only on some subset of $J^{1}E$. Indeed, it must be verified that the reduced space inherits a well-defined multicontact structure; whilst this was guaranteed for regular systems, it is not, a priori, clear that the same should be true in the singular case.
Having ascertained the necessary modifications of the current procedure to cases where degeneracies are present, our formalism is highly applicable to a broad array of contexts, several of which we shall now discuss. It has been demonstrated that, upon making a conformal decomposition $g_{\mu\nu}=e^{2\phi}h_{\mu\nu}$, with $\textrm{det}\,h_{\mu\nu}=-\,1$, of the spacetime metric, the Einstein-Hilbert Lagrangian admits a scaling symmetry, in which the scaling variable is precisely the conformal factor $\phi$ [24]. It will be of great interest to examine the details of this dynamical similarity in the first-order vielbein formalism [49, 50]. The use of the first-order construction is required so that we may work on the first jet bundle, which is multisymplectic; our motivation for employing the frame field and spin connection is that fermionic degrees of freedom may then be added, and the effects on the scaling symmetry examined.
In addition to General Relativity, many string-related low-energy effective actions are of a form in which a scaling symmetry is manifest. Consider, for example, the bosonic contribution to the NS-NS low-energy effective action of the type II theories [51]
$S_{\textrm{NS-NS}}=\frac{1}{2\kappa_{10}^{2}}\int d^{10}X\;\sqrt{-G}\,e^{-2\Phi}\left(\mathcal{R}-\frac{1}{2}|H_{3}|^{2}+4\,\partial_{\mu}\Phi\,\partial^{\mu}\Phi\right)$
The vector field $\partial_{\Phi}$ clearly generates rescalings of the action exactly of the type discussed throughout. Of course, more work is required to ascertain whether this is a true scaling symmetry; however, there is no obstruction to our calculating the forms $\theta_{L}^{\mu}$, and proceeding with the formal analysis.
In a similar spirit, there are many low-energy descriptions of non-Abelian gauge theories coupled to dilaton-like fields. As an example, actions of the form
$S=\int d^{4}x\,\left[-\frac{1}{4F(\Phi)}G_{\mu\nu}^{a}G^{\mu\nu}_{a}+\frac{1}{2}\partial_{\mu}\Phi\,\partial^{\mu}\Phi-V(\Phi)+J_{a}^{\mu}A^{a}_{\mu}\right]$
have been studied extensively in [52] and [53]. For certain forms of the function $F(\Phi)$, and non-perturbative dilaton potential $V(\Phi)$, this action possesses a scaling symmetry, in which we again identify $\Phi$ as the redundant degree of freedom. Of course, the examples provided are entirely qualitative in their analysis; however, they serve to appreciate that our construction is widely applicable to many theories of current interest in high-energy particle physics and relativistic field theory.
IX Conclusions and Outlook
We have advocated throughout that, while the canonical formulation has proved highly successful in the classical description, and subsequent quantisation of innumerable field theories, greater emphasis should be placed on dispensing with redundant mathematical structure. To this end, it was shown that, when described using the tools of multisymplectic geometry, many classical field theories possess degrees of freedom, corresponding to a global scale, whose presence is entirely superfluous, from the perspective of the dynamical evolution of the observables.
Following a brief review of the mathematical framework required to describe classical field theories in a manifestly covariant manner, we demonstrated that a multisymplectic Lagrangian (or Hamiltonian) system admitting a scaling symmetry could be reduced to a dynamically-equivalent description on a lower-dimensional space. This reduced theory was found to be action-dependent, and thus frictional in nature. Physically, we argued that this dissipative characteristic was to be interpreted as a necessary compensation in passing to a scale-insensitive description.
@@ IX Conclusions and Outlook
Through use of a pair of simple examples, we found that, in general, a theory of $N$ non-interacting massive scalar fields is dynamically-equivalent to a configuration of (N1)(N-1) massless fields moving in a non-zero potential, set by the value of the original field mass, together with an independent frictional component. Further, it was observed that any coupling between the scalar fields of the original theory destroys this separation into independent pieces, and introduces terms which mix the frictional degrees of freedom with those of the remaining fields.
From the results presented throughout, we have discussed a number of interesting areas of future work that may be explored. For example, it is clear that the most immediate requirement is a generalisation of our construction to singular theories, where the interaction between gauge and scaling symmetries is not fully understood. Within such a framework, it will of great interest to examine the solution space of a contact-reduced description of General Relativity, particularly at those points where the conventional formalism becomes singular and thus ill-defined. It is known, for instance, that within the cosmological sector, both FLRW and class-A Bianchi models may be extended beyond the initial singularity [54, 16].
@@
Progress has recently been made in the construction of a multisymplectic Hamiltonian description of a simple two-dimensional supersymmetric field theory [55]. Similarly, generalised Kähler geometries have been used to construct covariant Hamiltonian formulations of non-linear sigma models [56], and so an extension of our formalism to encompass both commuting and anticommuting variables appears tractable, and will undoubtedly prove enlightening.
@@
Finally, we have emphasised several times throughout the advantages presented by the covariant De-Donder Weyl formulation of classical field theories. Whilst a finite-dimensional phase space, and manifest covariance are attractive features of this approach, when contrasted with the canonical formalism, it cannot be overlooked that the latter provides a clear and rigorous path to quantisation, whereas the former does not [57]. In light of this, of particular interest is the development of a standardised framework, within which the restricted multimomentum bundle may be quantised, in much the same way that geometric quantisation offers a scheme for performing such calculations on the tangent bundle [58].